OpenStudy (anonymous):

whats the vertex and y intercept of the graph of the function y=-3(x+2)^2 +5

OpenStudy (anonymous):

This equation is in the form y = a(x - h)² + k where (h, k) is the vertex y = -3 (x + 2)^(2) + 5 y = -3 (x - -2)^(2) + 5 So the vertex is (-2, 5) The y-intercept is when x = 0 When x = 0, y = -3 * 4 + 5 = -7 So the y-intercept is (0, -7) And If you haven't done the vertex form of the equation just expand the equation so it's in the form y = ax² + bx + c so you get y = -3x² - 12x - 7 the x-coordinate of the vertex is given by x = -b/2a so x = 12/-6 = -2 And then subs x = -2 into the equation for y to get the y-coordinate of the vertex. The y-intercept is still found by putting x = 0 into the equation for y